# Protocol as Poetry: Case Study on Pak’s Protocol Arts

### v1

@@ Abstract.
Protocol art emerges at the confluence of blockchain-based smart contracts and a century-long lineage of conceptual art, participatory art, and algorithmic generative art practices. Yet existing definitions—most notably Primavera De Filippi’s ”protocolism”—struggle to demarcate this nascent genre from other art forms in practice. Addressing this definition-to-practice gap, this paper offers a focused case study of pioneering protocol artworks by Pak, an early and influential pseudonymous protocol artist who treats smart contracts as medium and protocol participation as message. Tracing the evolution from early open-edition releases of The Fungible‘ and the dynamic mechanics of Merge to the soul-bound messaging of Censored and the reflective absence of Not Found, we examine how Pak choreographs distributed agency across collectors and autonomous contracts, showing how programmable protocols become a social fabric in artistic meaning-making. Through thematic analysis of Pak’s works, we identify seven core characteristics that distinguish protocol art: (1) system-centric rather than object-centric composition, (2) autonomous governance for open-ended control, (3) distributed agency and communal authorship, (4) temporal dynamism and lifecycle aesthetics, (5) economic-driven engagement, (6) poetic message embedding in interaction rituals, and (7) interoperability enabling composability for emergence. We then discuss how these features set protocol art apart from adjacent artistic movements. By developing a theoretical framework grounded in Pak’s practice, we contribute to the emerging literature on protocolism while offering design implications for artists shaping this evolving art form.
Protocol Art, Smart Contract, Blockchain, Decentralized Control, System Art, Generative Art, Participatory Art, Open-endedness, Protocolism
$c$s: Human-centered computing Collaborative and social computing theory, concepts and paradigmsccs: Human-centered computing Interaction design theory, concepts and paradigms
Figure 1. The Logo of Pak. Pak’s symbol on Twitter is a circle known as ”The Nothing”.
@@ 1. Introduction
Over the last decade, the advent of blockchain networks, smart contracts, and non-fungible tokens (NFTs) has transformed digital culture by relocating trust from central institutions to cryptographic code: ”Code is Law” (Lessig, 2000). Beyond reshaping finance through DeFi, these technologies have begun to reconfigure artistic practice. An emerging cohort of artists now treats smart contracts as their medium, designing works whose meaning emerges through on-chain participation. Pioneering projects—terra0 (2016), which frames a forest as a self-owned entity  [^1] https://terra0.org/; Sarah Friend’s socially-driven on-chain Lifeforms (2021),  [^2] https://lifeforms.supply/; and the on-chain robo-botanical Plantoid (2017) series  [^3] https://plantoid.org/—demonstrate how smart-contract logic can serve as a medium to distribute creative agency through participation, while ”the aura of the artist is on maintaining the process” (Filippi, 2025). Recent exhibitions like the World Computer Sculpture Garden (2024)  [^4] https://worldcomputersculpture.garden/ foreground this shift, showcasing smart contracts as sculptural pieces that exist immutably on Ethereum. Together, these works signal the emergence of what is increasingly termed protocol art.
@@
Yet the concept of protocol art remains theoretically unsettled. In her recent article ”Protocolism,” De Filippi defines protocolism as an art form that ”uses protocols to guide diverse human and non-human executors in producing unique artworks,” where protocol means ”a set of rules or instructions that guide behaviour toward a specific outcome” (Filippi, 2025). While evocative, this definition is too broad in practice. It’s difficult to distinguish protocol art from other long-standing rule-based art forms: conceptual art, like Sol LeWitt’s Procedural Art (1976), Yoko Ono‘s ”Cut Piece” (1964), and Marina Abramović’s ”The Artist is Present” (2010); generative software art, such as Operator’s “Generative Choreography” (2023); and contemporary AI-assisted image models—all of which rely on rule-based logic. Even generative NFT platforms like Art Blocks invite collectors to mint algorithmic outputs but rarely cede autonomy to the code once deployed. not all runtime art, blockchain-native art, or on-chain art is protocol art, as maltefr claimed in her introductory essay to World Computer Sculpture Garden (maltefr, 2025). A gap persists between broad definitions of protocol art and the specific protocol art practices we observe in the emerging use of smart contracts as medium.
@@
Our research question is: What distinct characteristics of protocol art emerge from studying real-world protocol art practices?
To address this research question, we focus our investigation on the pseudonymous artist Pak, whose practice epitomizes protocol art’s aspirations by turning smart contracts into both medium and message. Active since 2020, Pak pioneered open-edition drops (e.g., The Fungible (2021)), dynamic NFTs that self-combine (e.g., Merge (2021)), and soul-bound tokens (Ohlhaver et al., 2022) that impose protest (e.g., Censored (2022)), extending most recently to the tokenized absence of #404 (2023). Each work demonstrates how participation through smart contracts can deliver poetic meaning: collectors are not merely owners but co-authors who experience the artwork through its evolving mechanics—whether fusion, auto-merging mass, or hidden blacklists. This participatory design transforms algorithmic logic into social reflection and has generated outsized cultural and economic impact; notably, Merge realized \$91.8 million in primary sales, then heralded as ”the largest art sale by a living creator” (Block, 2021). By embedding metaphors directly in code while enlisting thousands of participants, Pak offers a uniquely fertile corpus for examining the intertwined dynamics of protocol mechanics, collective engagement, and conceptual intent within emerging protocol art.
Using a qualitative grounded-theory approach, this study dissects Pak’s protocol artworks to theorise the behaviours they spark and delivers two key contributions. First, it presents a comparative matrix that evaluates Pak’s signature pieces through seven analytic lenses—concept, system-mechanism dynamics, participatory interaction, artist control, emergent behaviour, poetic meaning, and key statistics. Second, by rooting its analysis in the concrete mechanics of Pak’s oeuvre, the paper distils the emergent characteristics of protocol art, advances a clearer genre definition, maps its potential as an independent artistic movement, and outlines design implications for future creators.
Project (Year) Participatory Interaction System mechanism Artist control Participants Collective emergent behaviour
X (2020) 24-h open-edition mint on Nifty Gateway; collectors could mint unlimited copies within the window Time-scarcity → supply fixed by duration, not count None (rules immutable once deployed) Hundreds of editions; dozens of unique wallets FOMO-driven rush as deadline approached
The Title (2021) Fixed-price sale of visually identical NFTs, each bearing a different title/price Value experiment: image constant, metadata (title + price) varies Only pre-set prices/titles; tokens static One owner per piece (6 sold); headline Unsold (US\$1 M) remained unsold Community debate on where NFT value resides (image vs. narrative)
The Fungible (2021) Hybrid: open-edition “Cube” mint + puzzle/leaderboard rewards during Sotheby’s event Fungible cubes aggregate into higher-denomination NFTs; gameified rewards Rules and reward tiers fixed; artist validated puzzles + bonuses 23 598 cubes bought by 3 080 wallets → 6 156 NFTs total Competitive accumulation, cooperative puzzle-solving, viral social tasks
burn.art (2021) Anyone can burn any NFT → receives ASH ERC-20 via bonding curve “Creation via destruction”; burn rate halves every 5 M ASH Contract rules fixed, but artist sets future drops payable only in ASH Thousands of burns; thousands of ASH holders Emergent burn economy; strategising which NFTs to sacrifice
Lost Poets (2021) 65 536 “Page” NFTs sold; holders choose to burn for \$ASH or transform into AI-generated “Poet” NFT; on-chain renaming Multi-act game; burn-vs-mint branching; airdropped “Origin” Poets Pak triggers each Act and airdrops but cannot reverse user choices 8 k unique holders; 28 k Poets + remaining/burned Pages Strategy debates (keep, burn, name), puzzle-hunting, thematic naming guilds
Invisible Mechanism (2021) 30 critics airdropped non-transferable “Hate” NFTs Admin-only move() function; tokens “soul-bound” to artist Absolute; artist can seize/move tokens at will 30 forced recipients Wide discussion on ownership & power asymmetry in NFT contracts
Merge (2021) 48-h sale of “mass” units; every wallet holds a single NFT whose mass auto-adds when more bought or acquired Self-merging deflationary tokens; leaderboard tiers (Alpha, Beta …) Rules fixed; artist only designed incentives/ARG post-sale 28 983 buyers; 312 686 mass units → 28 984 NFTs Competitive hoarding, post-sale mergers reducing token count (“extinction game”); Internship DAO emerges to purchase ”Alpha”
Censored (2022) (a) 1/1 Clock auction (b) open-edition “Message” mint: any text, any price (a) Dynamic day-counter (b) Messages stay redacted; auto-reveal when/if Assange freed Clock autonomous; reveal switch bound to external event; artist cannot alter messages Clock owned by 10 000-member AssangeDAO; 30 000+ Message NFTs Grass-roots DAO fundraising; global participation, shared anticipation of reveal
[…]
Table 1. Pak’s Protocol Arts
@@ 2.1. Emerging Protocol Arts
Web3 technologies—blockchains, smart contracts, and non-fungible tokens (NFTs)—reconfigure the terms of authorship by shifting enforcement from institutions to code. Smart contracts are self-verifying, self-executing and tamper-resistant programs that guarantee rule compliance without a central server (Wang et al., 2018), while NFTs assign cryptographic provenance to digital artefacts. Together they allow artists to publish works that run autonomously and invite permissionless participation. As collectors interact on-chain —minting, trading, or triggering contract functions—the artwork’s state evolves in real time, making the protocol, rather than any single output, the true locus of creativity (Wu et al., 2025). Smart-contract infrastructure enables such participation at global scale. Web3 has revived—and radically extended—the tradition of participatory art: works in which the public co-produces the piece through prescribed actions.
Projects like terra0 (2016) imagine a forest that manages itself through a Decentralized autonomous organization (DAO), selling lumber rights to fund its own growth. Primavera De Filippi’s Plantoid (2017) embeds a robo-flower in Ethereum so that donors collectively “pollinate” the sculpture, triggering new commissions. The 2024 exhibition World Computer Sculpture Garden showcased artists—among them 0xhaiku, Absent, Loucas Braconnier, Sarah Friend, Material Protocol Arts, Rhea Myers et al., Paul Seidler—whose works exist solely as smart contracts, reinforcing the idea that the protocol itself can be material (maltefr, 2025). Most visibly, the pseudonymous artist Pak has released a string of high-profile blockchain works that invite thousands of collectors to fuse tokens, redact texts, or accumulate “mass,” demonstrating that participatory creativity can be both conceptual and commercially significant. Collectively, these examples point to an emerging genre—protocol art—in which rules encoded on-chain constitute the work, and audience transactions constitute its performance.
2.2. Defining Protocolism
Multiple disciplines offer definitions of a protocol. In computer networking, a protocol is ”a set of rules for formatting and processing data,” and network protocols function as ”a common language for computers” (Tanenbaum, 1981). De Filippi broadens this definition to ”a set of rules or instructions that guide behaviour toward a specific outcome” (Filippi, 2025). From the research initiative Summer of Protocols, dedicated to studying protocols as first-class concepts, Rao et al. describe it as ”a stratum of codified behaviour from which complex coordination can emerge” (Rao et al., [n. d.]). The material that implements a protocol may range from a written contract to cryptographically enforced software; what varies is the ”hardness of trust”—how difficult it is to deviate from the rules (Stark, 2023). The term protocolism has been advanced by De Filippi to capture practices that “use protocols to choreograph diverse human and non-human executors in the production of unique artworks” (Filippi, 2025).
Yet all these protoocl umbrella definition remains too broad when it faces art practice. Is Instruction-based conceptual works—Sol LeWitt’s Procedural Art (1976) an protocol art? Some of Sol LeWitt most famous pieces were not even executed by him in person. He did not sell paintings on canvases, instead he sold “recipes” or “procedures” (a.k.a “protocols”) with specific instructions that would then be implemented by a draftsman, who was required to faithfully execute the instructions, but whose own hand and judgement led to the final formal outcome of the work. For another example, Is the classic Participatory Art Marina Abramović - ”The Artist is Present” (2010) is a protocol art? Abramović invited viewers to sit across from her in a gallery while she maintained eye contact with them for extended periods, exploring themes of presence and connection. Modern example, do generative NFT platforms such as Art Blocks  [^5] https://www.artblocks.io/, where collectors mint outputs from pre-written algorithms. What, then, distinguishes protocol art?
Following the essay ”Computation in the Expanded Field,” (maltefr, 2025), maltefr argues that protocol art foregrounds the foundational structure that pre-configures a computational environment: the blockchain is not a mere distribution channel for images; it is the sculptural substrate itself. Protocol art is rooted in a century-long lineage of conceptual art, process art, participatory art, and algorithmic generative art practices. Yet it transcends these predecessors through its unique relationship with blockchain technology, smart contracts, permissionless participation, trustless infrastructure, and immutable landscape of affordances.
2.3. Pak and His Protocol Arts
@@
Pak— known as Murat Pak—is a pseudonymous crypto artist best known for harnessing blockchain smart-contracts as an artistic medium. After creating the AI-curation bot Archillect in 2014, Pak shifted to NFTs in 2020 and quickly became one of the field’s most commercially successful and conceptually adventurous figures: more than 180,000 of Pak’s tokens changed hands between 2020 and 2022, generating nearly US \$395 million in primary and secondary sales. The artist’s anonymity (some speculate “Pak” may be a collective) and code-driven practice challenge traditional ideas of authorship and ownership, while record-breaking projects such as Merge (US \$91.8 million) have placed Pak among the highest-grossing living creators in any medium.
For this case studies, we picked nine milestone protocol artworks from Pak: X (2020) introduced time-based scarcity by letting anyone mint unlimited editions for just 24 hours, shifting value from object to moment. The Title (2021) sold the same image under differently priced names, proving that metadata—rather than pixels—can dictate market value. The Fungible (2021) used a Sotheby’s open-edition drop to turn fungibility itself into spectacle, releasing thousands of interchangeable tokens that questioned uniqueness. burn.art / \$ASH (2021) created a perpetual “creation-through-destruction” loop in which collectors burn any NFT to mint the social currency ASH, gamifying scarcity. Lost Poets (2021) unfolded as a multi-stage strategy game of 65,536 AI-generated “Pages” that could evolve, merge, or be burned, turning collectors into co-writers of an ever-shifting narrative. Invisible Mechanism (2021) premiered the “move” contract, giving the artist the unilateral power to relocate tokens—an audacious commentary on creator sovereignty hidden beneath a drop nicknamed Hate. Merge (2021) sold 312,686 “mass units” that algorithmically fused into larger single tokens after every transfer, making ownership a live sculptural process. Censored (2022), made with Julian Assange, paired a DAO-funded one-of-one “Clock” counting Assange’s days in prison with open-edition “Censored” tokens, weaponising provenance for political protest. Not Found #404 (2023) assigned token ID 404 to an NFT deliberately devoid of metadata, transforming the web’s “page not found” error into a meditation on digital absence.

### v2

# Protocol as Poetry: A Case Study of Pak’s Smart Contract-Based Protocol Art

@@ Abstract.
Protocol art has recently proliferated through blockchain-based smart contracts, building on a century-long lineage of conceptual, participatory, interactive, systematic, algorithmic, and generative art practices. Few studies have examined the characteristics and appreciation of this emerging art form. To address this gap, this paper presents an annotated portfolio analysis of protocol artworks by Pak, a pioneering and influential pseudonymous artist who treats smart contracts as medium and collective participation through protocol as message. Tracing the evolution from early open-edition releases of The Fungible (2021) and the dynamic mechanics of Merge (2021) to the soul-bound messaging of Censored (2022) and the reflective absence of Not Found (2023), we examine how Pak choreographs distributed agency across collectors and autonomous code, demonstrating how programmable protocols become a social fabric in artistic meaning-making. Through thematic analysis of Pak’s works, we identify seven core characteristics distinguishing protocol art from other art forms: (1) system-centric rather than object-centric composition, (2) autonomous governance enabling open-ended control, (3) distributed agency and communal authorship, (4) temporal dynamism and lifecycle aesthetics, (5) economy-driven engagement, (6) poetic message embedded in interaction rituals, and (7) interoperability enabling composability for emergent complexity. We then discuss how these features set protocol art apart from adjacent movements such as conceptual, generative, participatory, interactive, and performance art. By analyzing principles grounded in Pak’s practice, we contribute to the emerging literature on protocol art (or “protocolism”) and offer design implications for future artists exploring this evolving form.
Protocol Art, Smart Contract, Blockchain, Decentralized Control, System Art, Generative Art, Participatory Art, Open-endedness, Protocolism, Systems Aesthetics, Collective Behavior
journalyear: 2025copyright: ccconference: 12th International Conference on Digital and Interactive Arts : ARTECH 2025 – Media Art Cultures, Communities & Territories; November 26–28, 2025; Braga, Portugalbooktitle: 12th International Conference on Digital and Interactive Arts : ARTECH 2025 – Media Art Cultures, Communities & Territories (ARTECH 2025), November 26–28, 2025, Braga, Portugaldoi: 10.1145/3773699.3773918isbn: 979-8-4007-2001-7/25/11ccs: Human-centered computing Collaborative and social computing theory, concepts and paradigmsccs: Human-centered computing Interaction design theory, concepts and paradigmsccs: Applied computing Media artsccs: Information systems Multimedia information systems
Refer to caption A white torus (ring shape) centered on a solid black background, serving as the recurring visual identity of pseudonymous digital artist Pak.
Figure 1. Pak’s signature symbol, a minimalist white ring on black known as “The Nothing,” used as the artist’s persistent identity across social media and NFT platforms.
@@ 1. Introduction
Over the last decade, the advent of blockchain networks, smart contracts, and non-fungible tokens (NFTs) has transformed digital culture by relocating trust from central institutions to cryptographic code—famously encapsulated in the motto “code is law” (Lessig, 2000). Beyond reshaping finance through decentralized finance (DeFi), these technologies have begun to reconfigure artistic practice. An emerging cohort of artists now treat smart contracts as their medium, designing works whose meaning unfolds through on-chain participation. Pioneering projects such as Paul Seidler’s terra0 (2016) [^1] https://terra0.org/ (Paul Seidler, 2016), which frames a forest as a self-owning economic entity (Seidler et al., 2017), Sarah Friend’s socially-driven on-chain Lifeforms (2021) [^2] https://lifeforms.supply/ (Friend, 2021), and Primavera De Filippi’s robo-botanical Plantoid series (2015) [^3] https://plantoid.org/ (Filippi, 2015) illustrate how smart-contract logic can serve as a medium to distribute creative agency through participation. In these works, “the aura of the artist is on maintaining the process” (De Filippi, 2024), as code and crowd together generate the artistic outcome. Recent exhibitions such as World Computer Sculpture Garden (2024) [^4] https://worldcomputersculpture.garden/ (0xhaiku et al., 2024) have further foregrounded this shift, showcasing smart contracts as sculptural pieces that exist immutably on Ethereum: “This permissionless computing platform has been reimagined as one, functioning simultaneously as archive, gallery, and economy.” (Rauch, 2024) Together, these examples signal the emergence of what is increasingly termed protocol art.
@@
Yet the concept of protocol art remains theoretically unsettled. De Filippi (De Filippi, 2024) defines protocol art as an art form where “protocols serve as the bridge between two pivotal figures: the original artist (i.e., the ‘protocol artist’) and a diverse array of executors (both human and non-human).” These protocols “comprise sets of instructions and constraints meticulously crafted to guide the production of unique artworks” (De Filippi, 2024). While evocative, this definition is very broad in practice. It is difficult to distinguish protocol art from other long-standing rule-based art forms. For example, instruction-based conceptual art—such as Sol LeWitt’s (LeWitt, 1967) procedural Wall Drawing series (MASS MoCA, 2021), Yoko Ono’s participatory Cut Piece (1964) (Ono, 1964), or Marina Abramović’s durational The Artist Is Present (2010) (Biesenbach, 2010)—relies on rules and audience action. Generative software artworks follow similar principles, including Vera Molnár’s early algorithmic art and contemporary AI-generated pieces (Molnar, 1975). Recent blockchain-based generative NFT platforms such as Art Blocks [^5] https://artblocks.io/ invite collectors to mint their own algorithmic editions—originally generated from art code in smart contracts. Although some pieces are interactive, allowing interaction through webpages, they rarely involve dynamic interaction and evolution at the smart-contract level once minted. “The blockchain is not a distribution channel for images but rather the material of the art itself.” as curator Malte Rauch (@maltefr) argued in his introductory essay to World Computer Sculpture Garden (Rauch, 2024). A gap persists between broad definitions of protocol art and the specific protocol-art practices now emerging from the use of smart contracts as artistic medium.
@@
Our research question is: What distinct characteristics define protocol art as evidenced through real-world practices?
We address this gap through an annotated portfolio analysis (Gaver and Bowers, 2012; Culén et al., 2020) of the pseudonymous artist Pak [^6] https://x.com/muratpak, who is widely credited with establishing protocol art’s foundational vocabulary. Our focus on a single artist is methodologically purposive rather than a sampling limitation: in emerging fields, paradigmatic cases—exemplars that establish the conventions and possibilities of a new form—merit intensive single-case analysis. Just as Duchamp’s readymades warranted deep study to understand conceptual art’s foundations, or as John Cage’s compositions required close analysis to theorize indeterminacy in music, Pak’s protocol artworks represent the paradigmatic instantiation of smart-contract-as-medium practice. Pak did not merely participate in protocol art—Pak effectively invented its vocabulary: open-edition time-scarcity (X, 2020), deflationary merging mechanics (Merge, 2021), burn-to-mint economies (burn.art, 2021), soul-bound protest tokens (Censored, 2022) (Weyl et al., 2022), and tokenized absence (Not Found, 2023). These mechanisms have since been widely imitated across thousands of NFT projects but never surpassed in conceptual sophistication. Each work demonstrates how participatory on-chain interaction can deliver poetic meaning: collectors are not merely purchasers but co-authors who experience the artwork through its evolving mechanics—whether through token fusion, auto-merging mass, or hidden messages revealed only under certain conditions. This participatory design transforms algorithmic logic into a collective movement that prompts social reflection and generates significant cultural and economic impact. Notably, Merge realized \$91.8 million in primary sales in December 2021, heralded at the time as “the largest-ever public sale of an artwork by a living artist” (Block, 2021; Thomas, 2021). By embedding metaphors directly in code while enlisting thousands of participants, Pak’s oeuvre offers a fertile—and uniquely well-documented—corpus for examining the intertwined dynamics of protocol mechanics, collective engagement, and conceptual intent.
In the remainder of this paper, we analyze Pak’s key works to derive common characteristics of protocol art. We present seven salient characteristics that recur across Pak’s protocol artworks, which together form what we term Pak’s protocol aesthetic. Then, we discuss how these traits distinguish protocol art from adjacent art movements such as conceptual, systemic, generative, participatory, interactive, and performance art, and we outline limitations of our single-artist study along with directions for future research. We conclude by reflecting on how protocol art (or “protocolism”) expands the discourse of digital art and offers new opportunities for artists to engage code, community, and concept as one autonomous system.
@@ 2.1. Emerging Protocol Arts
Web3 technologies—blockchains, smart contracts, and non-fungible tokens (NFTs)—reconfigure the terms of authorship by shifting enforcement from institutions to code (Sidorova, 2025). Smart contracts are self-verifying, self-executing, tamper-resistant programs that guarantee rule compliance without a central authority (Wang et al., 2018). NFTs assign cryptographic provenance and ownership to digital artifacts. Together, these technologies allow artists to publish works that run autonomously and invite permissionless participation. As collectors interact on-chain—minting, trading, or triggering contract functions—the artwork’s state can evolve in real time, making the protocol, rather than any single static output, the true locus of creativity (Wu et al., 2025; Guan and others, 2025). Smart-contract infrastructure enables this participatory evolution at global scale. Chris Dixon argues in “Read Write Own” (Dixon, 2024) that ownership distinguishes Web3 from Web2’s write and Web1’s read—blockchain networks grant power and economic benefits to communities of users. By reshaping ownership, web3 has revived and radically extended the lineage of participatory art (Bishop, 2023): art in which the public co-owns and co-produces the piece through autonomous participation governed by protocols.
Several early projects foreshadowed the protocol art paradigm. terra0 (2016) presented a vision of an artwork that manages itself: a forest equipped with sensors and coded as a Decentralized Autonomous Organization (DAO) that sells its lumber rights to fund its own growth (Paul Seidler, 2016; Seidler, 2025). Primavera De Filippi’s Plantoid (first created in 2015) (Filippi, 2015) embeds a robotic “flower” sculpture in Ethereum such that donors collectively “pollinate” the piece with cryptocurrency to fund the creation of new offspring sculptures. By 2024, exhibitions such as World Computer Sculpture Garden (0xhaiku et al., 2024) showcased artists (among them 0xhaiku, Absent, Loucas Braconnier (Figure31), Sarah Friend, Material Protocol Arts, Rhea Myers et al., Paul Seidler) whose works exist solely as on-chain contracts, reinforcing the idea that the protocol itself can be the artistic material (Rauch, 2024). Rhea Myers’s collected writings further document how smart contracts have been used as an artistic medium since 2011 (Myers, 2023).
@@
The pseudonymous artist Pak—also known as Murat Pak—is best known for harnessing blockchain smart contracts as an artistic medium. Early work included creating the AI-curation bot Archillect [^7] https://archillect.com/ in 2014. Pak shifted to smart contracts in 2020 and quickly became one of the field’s most commercially successful and conceptually adventurous figures. The artist’s anonymity (some speculate “Pak” may be a collective) and code-driven practice challenge traditional ideas of authorship and ownership. Record-breaking projects such as Merge have placed Pak among the highest-grossing living creators in any medium (Thomas, 2021). Pak’s projects innovate in protocol mechanisms, inviting thousands of collectors to fuse tokens, burn assets for new tokens, redact and reveal messages, or accumulate “mass”—demonstrating that participatory creativity can be both conceptually rich and commercially significant.
Collectively, these examples point to an emerging genre—protocol art—in which the rules encoded on-chain constitute the work, and the audience’s transactions constitute its performance.
2.2. Defining Protocol Art?
Multiple disciplines offer definitions of protocol. In computer networking, a protocol is “a set of rules for formatting and processing data,” a common language that enables coordination among machines (Tanenbaum, 1981). Media theorist Alex Galloway famously broadened this concept: “protocol” is the principle of organization that governs distributed systems—“how control exists after decentralization” (Galloway, 2001). Venkatesh Rao and colleagues, in the Summer of Protocols [^8] https://summerofprotocols.com/ research initiative, describe protocols as “a stratum of codified behavior from which complex coordination can emerge” (Rao et al., 2024). The material embodiment of a protocol can range from written rules to cryptographically enforced software; what varies is the “hardness of trust”—i.e. how difficult it is to deviate from the rules (Stark, 2022).
Within the arts, in her essay “Protocolism: The Evolving Landscape of Art in the Age of AI” (De Filippi, 2024), originally published in All Media is Training Data, protocolism has been advanced by De Filippi to describe practices that use protocols to choreograph diverse human and non-human executors in the production of artworks (De Filippi, 2024). However, as noted above, this umbrella definition can be too broad when applied to art. Are instruction-based conceptual works like Sol LeWitt’s wall drawings “protocol art”? One might argue yes: LeWitt often did not execute his own wall drawings—he sold the “recipe” or instructions (a kind of artistic protocol) to be faithfully executed by others, meaning the artwork is essentially a set of rules carried out by participants. Similarly, is a classic participatory performance piece like Marina Abramović’s The Artist Is Present (2010) (Biesenbach, 2010) a protocol artwork? The piece had a simple rule: viewers took turns sitting silently across from the artist for a duration, exploring presence and connection. One could view this as a human protocol enacted in a gallery. Even modern generative NFT platforms such as Art Blocks, where collectors mint outputs from pre-written algorithms, involve protocols of creation. What, then, distinguishes protocol art in the context of Web3?
Following the essay “Computation in the Expanded Field” (Rauch, 2024), Malte Rauch argues that true protocol artists “are primarily concerned with foundational structures or sets of rules that precede and define a computing environment”. In these works, “The blockchain is not a distribution channel for images but rather the material of the art itself.” The artistic focus shifts from static outputs to the system of rules (Rivero-Moreno, 2024; Poposki, 2024b). In other words, what sets protocol art apart is that the rule-set autonomously generates and governs the art experience over time, without further intervention. Traditional conceptual or generative art (Boden and Edmonds, 2009) uses instructions or code to produce a finished piece—perhaps an interactive one. Protocol art, however, uses instructions or code to create an open-ended process or system that invites collective participation through co-creation of meaning and co-ownership of the piece. As we will show through Pak’s work, protocol art emphasizes system-centric creation, persistent autonomy, communal participation, temporal evolution, integrated economic and social incentives, and cross-project interoperability—together resulting in a distinctive aesthetic and conceptual stance. However, the terrain of protocol art continues to be shaped as more artists join this movement, with emerging critical (Poposki, 2025, 2024a) and ethical (Calvo, 2024) perspectives enriching the discourse. Defining protocol art is still far from consensus and remains an ongoing process.